### v1

# Efficient Reasoning for LLMs through Speculative Chain-of-Thought

Abstract
Large reasoning language models such as OpenAI-o1 and Deepseek-R1 have recently attracted widespread attention due to their impressive task-solving abilities. However, the enormous model size and the generation of lengthy thought chains introduce significant reasoning costs and response latency. Existing methods for efficient reasoning mainly focus on reducing the number of model parameters or shortening the chain-of-thought length. In this paper, we introduce Speculative Chain-of-Thought (SCoT), which reduces reasoning latency from another perspective by accelerated average reasoning speed through large and small model collaboration. SCoT conducts thought-level drafting using a lightweight draft model. Then it selects the best CoT draft and corrects the error cases with the target model. The proposed thinking behavior alignment improves the efficiency of drafting and the draft selection strategy maintains the prediction accuracy for complex problems. Experimental results on GSM8K, MATH, GaoKao, CollegeMath and Olympiad datasets show that SCoT reduces reasoning latency by 48$\sim$66for Deepseek-R1-Distill-Qwen-32B while achieving near-target-model-level performance. Our code is available at https://github.com/Jikai0Wang/SpeculativeCoT.
1 Introduction
Large-scale reasoning language models (e.g., OpenAI-o1 (Jaech et al.,, 2024) and Deepseek-R1 (DeepSeek-AI et al.,, 2025)) have exhibited exceptional performance across a broad range of reasoning scenarios. Their superior performance stems from two key factors: massive model scale (e.g., DeepSeek-R1 with 671B parameters) and long chain-of-thought (CoT) (Wei et al.,, 2022) generation (reaching over 10,000 tokens for complex tasks). However, both factors incur substantial computational overhead, resulting in huge inference costs. In many reasoning systems where the model’s reasoning process remains implicit, excessive reasoning time induces significant response latency.
Recent works (Paliotta et al.,, 2025; Liao et al.,, 2025; Xu et al., 2025b, ; Zhang et al.,, 2025) mainly tackle this challenge by developing smaller reasoning models or reducing the generation length of CoTs. Distillation (Paliotta et al.,, 2025; Liao et al.,, 2025) serves as a commonly used technique for transferring strong reasoning capabilities from large reasoning models to small ones. Representatively, Deepseek-R1-Distill-Qwen-1.5B (DeepSeek-AI et al.,, 2025) exhibits robust performance on diverse datasets. Approaches such as StepSkip (Liu et al.,, 2024) and TokenSkip (Xia et al.,, 2025) suggest training reasoning models on compressed CoT data to shorten the CoT length during inference. In practical applications, reasoning models need to solve problems with widely varying complexity levels. Fundamentally, employing large reasoning models to handle simple tasks is a waste. On the other hand, using small models alone or compressing the chain-of-thought length may compromise performance on complex tasks. Now that we have a powerful and efficient small model, why not prioritize this lightweight alternative and only invoke the large model for deeper thinking when the problem exceeds the capability threshold of the small one?
Figure 1: Average reasoning accuracy and efficiency of Speculative Chain-of-Thought (SCoT) with Deepseek-R1-Distill-Qwen-32B and Deepseek-R1-Distill-Qwen-1.5B on GSM8K, MATH, GaoKao, CollegeMath and Olympiad. Graph (a): SCoT achieves near-target-model performance level on the 5 datasets. Graph (b): SCoT demonstrates an average speed-up ratio of 2.3× for reasoning latency. Graph (c): Comparison of reasoning efficiency-accuracy trade-off.
Based on this motivation, we propose Speculative Chain-of-Thought (SCoT), a reasoning framework that reduces reasoning latency through the collaboration of large and small models. The core idea of SCoT is to use a lightweight model to efficiently generate multiple CoT drafts which are then picked and used by the target model. SCoT is inspired by speculative decoding (Stern et al.,, 2018; Xia et al.,, 2023; Leviathan et al.,, 2023), which performs token-level speculation, whereas our approach conducts thought-level speculation. Unlike shortening the length of CoT, SCoT accelerates the thinking process from another dimension, that is, by increasing the speed of CoT generation.
In Section 3, we introduce thinking behavior alignment to improve drafting efficiency. We also design an error correction mechanism to preserve accuracy in dealing with complex tasks. We construct a training set by 500 samples from GSM8K for fine-tuning the target model to improve the capability of draft selection and error detection. We conduct comprehensive experiments in Section 4 to evaluate the effectiveness of SCoT. Figure 1 displays the overview reasoning accuracy and efficiency of SCoT. Adapting to task difficulty and draft quality, the draft selection mechanism decides when the target model should rethink. Simple tasks mainly utilize the draft model for thought chain generation, while complex problems engage the target model more extensively in the reasoning process. SCoT reduces reasoning latency by 48$\sim$66on 5 datasets (including the Olympiad-level dataset (He et al.,, 2024)) while maintaining near-target-model-level performance.
2 Related Works
Efficient Reasoning Reasoning LLMs (Jaech et al.,, 2024; DeepSeek-AI et al.,, 2025) generate lengthy CoTs to enhance performance, albeit at the cost of increased latency. Some approaches (Ding et al.,, 2024; Luo et al.,, 2025; Xia et al.,, 2025; Xu et al., 2025b, ; Zhang et al.,, 2025) speed up the thinking process by reducing the length of the CoT at the token level. Liu et al., (2024) and Xia et al., (2025) use compressed CoTs to fine-tune the reasoning LMs. Hou et al., (2025) prunes long CoTs through reinforcement learning. Xu et al., 2025a and Aytes et al., (2025) guide the model to think efficiently through prompting. Soft CoT methods (Saunshi et al.,, 2025; Xu et al., 2025b, ; Hao et al.,, 2024) use latent space representations to replace reasoning tokens, thus reducing the excessive redundancy in the reasoning process. Another mainstream approach (Yu et al.,, 2024; Paliotta et al.,, 2025; Liao et al.,, 2025) for efficient reasoning is to teach smaller models to think. Paliotta et al., (2025) and Liao et al., (2025) transfer reasoning ability from large models to small ones by distillation. Liu et al., (2025) study the impact of quantization for reasoning LLMs. Our approach leverages collaborative inference between large and small models, combining the small model’s efficiency with the large model’s accuracy to achieve an improved efficiency-accuracy trade-off in reasoning LLMs.
Speculative Decoding Autoregressive language models generate one token in one decoding step, which greatly limits its decoding efficiency. Speculative decoding (Stern et al.,, 2018; Xia et al.,, 2023; Leviathan et al.,, 2023; Chen et al., 2023a, ) adopts lightweight models for drafting and target models for parallel verification to generate multiple tokens in one decoding step, thus achieving lossless acceleration. The draft model can be an independent small model (Leviathan et al.,, 2023; Chen et al., 2023a, ; Spector and Re,, 2023; Chen et al., 2023b, ), a submodule of the target model (Cai et al.,, 2024; Zhang et al.,, 2023; Li et al.,, 2024), or an external corpus (He et al.,, 2023). Draft tokens that match the original output of the target model will be accepted under a strict verification strategy. Most speculative decoding methods perform token-level drafting, while we borrow the idea of speculation to perform thought-level drafting to improve the reasoning efficiency of LLMs.
3 Speculative CoT
This section introduces Speculative Chain-of-Thought (SCoT), an algorithm for accelerating LLM reasoning. SCoT employs a small model to draft reasoning chains while utilizing the target model to select the best draft for generating the final answer. The overview of SCoT is shown in Figure 3.
3.1 Formulation
For question $q$ in the given question set Q, the reasoning model $M$ first enters the thinking stage after inputting $q$. Specifically, M will output a chain-of-thought $T=(y_{\_}{1},y_{\_}{2},...,y_{\_}{l})$ starting and ending with pre-trained special tokens, where $y_{\_}{i}$ represents the $i$-th output token and $l$ is length of $T$. Take Deepseek-R1 (DeepSeek-AI et al.,, 2025) as an example. The thinking process starts with “<think>” and ends with “</think>”. Next, the model generates the final answer $A$ based on $q$ and $T$. Our goal is to reduce the inference overhead of $T$ while ensuring the quality of the final answer.
3.2 Generating Drafts of Thoughts
Figure 2: Comparison of reasoning accuracy on GSM8K.
First, we start with a question: “Can small models generate high-quality chain-of-thoughts?”. We sampled 500 samples in GSM8K (Cobbe et al.,, 2021) to evaluate the reasoning accuracy of Deepseek-R1-Distill-Qwen-1.5B and Deepseek-R1-Distill-Qwen-32B (DeepSeek-AI et al.,, 2025). Figure 2 displays the results. Under the greedy sampling method, the performance of the 1.5B model is significantly worse than that of the 32B model. However, if we allow the 1.5B model to provide 5 answers under the nuclear sampling method and select the best one among the 5 replies for each sample, its overall accuracy on GSM8K is close to that of the 32B model.
From the above experiments, we can conclude that the small model has the potential to generate high-quality thought chains, but a stronger verifier is needed to control or select its generated results. Therefore, in order to accelerate the thinking process of reasoning LLMs, we first use a smaller model $M_{\_}{d}$ to draft $n$ chains of thoughts in parallel:
[…]
We use nuclear sampling for the small model thinking to enhance the diversity of drafts of CoTs. Thanks to their smaller parameter size, smaller models can draft reasoning chains significantly faster than large models when the length of the CoT is the same, thus speeding up the thinking process.
Figure 3: Overview of Speculative Chain-of-Thought. Given a question $Q$, SCoT first applies a lightweight draft model to generate multiple CoT drafts in parallel. The draft model is fine-tuned with LoRA modules to align the thinking behavior of the target model. It appends a special CoT option for the case where all drafts are wrong. The target model for selecting the best CoT draft is also fine-tuned with LoRA modules for improved accuracy. With the designed prompt template $S$, only one forward propagation is needed to get the index of the best CoT. Once the best CoT is picked, SCoT directly adopts it for the original target model to generate the final answer. Only if no draft is selected, SCoT will rethink with the target model to ensure the quality of the generated answer.
3.3 Thinking Behavior Alignment
Although small models have higher throughput than large ones, significant differences in their reasoning behavior may still result in limited, or even reduced, reasoning speed. In our experiments, we observed that even models of the same series can exhibit significant differences in thinking behavior. The inference speed of Deepseek-R1-Distill-Qwen-1.5B is about 6 times that of Deepseek-R1-Distill-Qwen-32B. However, the average CoT length of the 1.5B model on the GSM8K dataset is more than 4 times that of the 32B model. The CoT of the 1.5B model contains more redundant tokens and repeated reflections. Therefore, the draft model requires training in alignment of thinking behavior to ensure drafting efficiency.
We adopt 1500 samples from the GSM8K training set to make the target model generate the corresponding thought chains. The training data $d=(x_{\_}{1},x_{\_}{2},...,x_{\_}{m},y_{\_}{1},y_{\_}{2},...,y_{\_}{l})$ is a combination of the question and generated CoT, where $m$ is the length of the question and $x_{\_}{\leq m}$ are input tokens. Then we use LoRA modules (Hu et al.,, 2022) to fine-tune the draft model with the cross-entropy loss:
_$\mathcal{L}_{\_}{Draft}=-\frac{1}{l}\sum_{\_}{i=1}^{l}\log p_{\_}{M_{\_}{d}}(y_{\_}{i}|x_{\_}{\leq m},y_{\_}{<i}),$ (2)
where $p_{\_}{M_{\_}{d}}(\cdot|\cdot)$ represents the output probability distribution of $M_{\_}{d}$. The original parameters of the draft model are frozen during training. The LoRA modules are adapted to the $\boldsymbol{Q}$ and $\boldsymbol{V}$ matrices in the attention block of each layer. After training, we merge the LoRA module parameters into the weight matrices of the draft model. Therefore, there is no additional overhead during inference.
3.4 Draft Selection and Error Correction
After generating CoT drafts ($T_{\_}{1},T_{\_}{2},...,T_{\_}{n}$) by the fine-tuned draft model, we design a prompt template $S$ for selecting these thought chains. Figure 3 displays the specific content of this template. For particularly challenging problems where all CoT drafts may be incorrect, we aim for the target model to detect such cases to ensure the accuracy of the final answer. For this purpose, we introduce a special option $T_{\_}{n+1}$, which indicates that all CoT drafts are wrong. Through a single forward propagation, we can efficiently obtain the CoT index with the highest probability from the output distribution of the next token by the target model:
$index=\underset{i\in\mathcal{V}}{\arg\max}\,P_{\_}{M}(i|S(q,T_{\_}{\leq n+1})),\,\mathcal{V}=\{1,2,...,n+1\}.$ (3)
To improve the accuracy of the target model in selecting the correct CoT drafts and detecting errors, we also use the LoRA modules to fine-tune the target model. We use 500 samples from the GSM8K training set to construct the training data. For each sample, we use the draft model to generate $n$ thought chains and deploy template $S$ as the input. We judge the correctness of the CoT drafts based on the ground truth and construct the label set $\mathcal{Y}$, which is the set of indices for all correct drafts. See Appendix A for examples. We design the following loss function to fine-tune the target model:
_$\mathcal{L}_{\_}{Target}=\min\{-\log p_{\_}{M}(y|S(q,T_{\_}{\leq n+1}))|y\in\mathcal{Y}\}.$ (4)
Likewise, the original model parameters are frozen, and parameters in LoRA modules are merged into the weight matrices of $\boldsymbol{Q}$ and $\boldsymbol{V}$. The fine-tuned target model is only used for CoT selection.
There are two situations when selecting CoT. The first is that the output $index<n+1$, that is, the target model selects the best one $T_{\_}{index}$ among the CoTs generated by the draft model. In this case, we directly let the target model generate the final answer based on $q$ and $T_{\_}{index}$. The second situation is the output $index=n+1$, which indicates that the problem is too hard and the draft model failed to produce correct reasoning. To maintain performance in the face of these complex problems, we instruct the target model to rethink the question and generate the final answer.
4 Experiments
4.1 Settings
We adopt Deepseek-R1-Distill-Qwen-32B (DeepSeek-AI et al.,, 2025) (abbreviated as D-Qwen-32B) as the target model and Deepseek-R1-Distill-Qwen-1.5B as the draft model. We use a single A100-PCIE-40GB GPU for the 1.5B model and 4 A100-PCIE-40GB GPUs for the 32B model. We evaluate the efficiency and accuracy of Speculative CoT on reasoning datasets of various difficulty levels, including GSM8K (Cobbe et al.,, 2021), MATH (Hendrycks et al.,, 2021), GaoKao-En-2023 (Liao et al.,, 2024), CollegeMath (Tang et al.,, 2024) and Olympiad (He et al.,, 2024). As stated in Section 3, we use data from GSM8K training set to fine-tune the models with LoRA modules. Note that there is no overlap in the data used for training and evaluation. We set the rank of LoRA modules to 8 for both the draft and target model. The number of CoT drafts $n$ is set to 5. To avoid running out of memory when selecting the drafts, the maximum draft length is set to 5000. We compare the reasoning latency and accuracy of Speculative CoT with the vanilla reasoning by the 32B model.
4.2 Main Results
Table 1: Comparison of CoT length and reasoning efficiency. We report average CoT length of the target model $l_{\_}{M}$, average CoT length of the draft model $l_{\_}{M_{\_}{d}}$ and average CoT latency of each sample $t(s)$. Speed-up ratio $r$ is calculated by average thinking time compared with vanilla decoding.
Method GSM8K MATH GaoKao CollegeMath Olympiad
$l_{\_}{M}$ $l_{\_}{M_{\_}{d}}$ $l_{\_}{M}$ $l_{\_}{M_{\_}{d}}$ $l_{\_}{M}$ $l_{\_}{M_{\_}{d}}$ $l_{\_}{M}$ $l_{\_}{M_{\_}{d}}$ $l_{\_}{M}$ $l_{\_}{M_{\_}{d}}$
D-Qwen-32B 302 - 2123 - 3230 - 1051 - 8431 -
SCoT(32B+1.5B) 45 265 151 1855 630 2767 110 1136 3553 4194
$t$ $r$ $t$ $r$ $t$ $r$ $t$ $r$ $t$ $r$
D-Qwen-32B 26.5 1.00 225.1 1.00 369.5 1.00 95.3 1.00 1092.2 1.00
SCoT(32B+1.5B) 11.8 2.26 77.0 2.92 160.1 2.31 43.7 2.18 575.7 1.90
Lower Reasoning Latency  Table 1 demonstrates the average CoT length and the average CoT latency of each sample on the 5 datasets. $l_{\_}{M}$ and $l_{\_}{M_{\_}{d}}$ represent the numbers of tokens generated by the target model and the draft model during the thinking stage, respectively. Since the draft model generates 5 draft chains in parallel, $l_{\_}{M_{\_}{d}}$ refers to the length of the longest draft CoT. The average latency $t$ is measured in seconds. For the 32B model, $l_{\_}{M}$ exhibits a wide range of 302 to 8431 across different datasets. Except for MATH where the average CoT length of SCoT ($l_{\_}{M}+l_{\_}{M_{\_}{d}}$) is shorter than that of the 32B model, the overall CoT lengths of the two methods are close. For SCoT, most of the reasoning tokens are generated by the draft model, and the CoT tokens generated by the target model account for 8.8$\sim$45.9of the total CoT tokens. As task difficulty increases, small models exhibit higher error rates, thus necessitating greater involvement from the target model in the reasoning process. The draft selection mechanism dynamically determines whether the target model should reconsider its output, evaluating both problem complexity and draft generation quality. For simpler tasks, thought chains are predominantly produced by the draft model. Conversely, the system allocates more thought chain generation to the target model when handling complex problems.
The throughput of the 1.5B model is about 5 times that of the 32B model in our experimental environment. The CoT latency of SCoT is mainly composed of the target model reasoning time, draft model reasoning time, and draft selection time. For the GSM8K dataset, the proportions are approximately 33.9%, 60.2%, and 5.9% respectively. Generally, SCoT achieves greater speedup ratios on simpler tasks where the CoT drafts demonstrate higher quality. SCoT achieves a speedup ratio ranging from 1.90× to 2.92× compared to vanilla reasoning.
Near-target-model Level Accuracy  Figure 4 displays the accuracy of SCoT on these datasets. Comparative results indicate that SCoT maintains near-target-model performance levels on all datasets (even for the Olympiad-level task), with clearly superior accuracy compared to standalone draft model reasoning. For CollegeMath, SCoT even achieves a lossless acceleration.
Figure 4: Comparison of accuracy of the final answer on the 5 datasets.
4.3 Ablation Studies
Table 2: Comparative analysis of average CoT draft length on DeepSeek-R1-Distill-Qwen-1.5B with versus without thinking behavior alignment training.
Method GSM8K MATH GaoKao CollegeMath Olympiad
Deepseek-R1-Distill-Qwen-1.5B 606 2092 2589 1355 4032
+ Thinking Behavior Alignment 193 1183 1487 709 3554
We conduct an ablation study in this section. We carry out 3 sets of experiments to observe the changes in accuracy on the 5 datasets with Deepseek-R1-Distill-Qwen: 1) SCoT with Single Draft: Only one CoT draft is generated for each sample (except for the special draft $T_{\_}{n+1}$), 2) SCoT without Tuning $M$: We use the original target model instead of the fine-tuned version to select the drafts, 3) SCoT without Error Correction: We remove $T_{\_}{n+1}$ and the rethinking strategy. Table LABEL:tab:ma shows the results. SCoT exhibits the highest accuracy on all datasets, indicating that both the three proposed techniques contribute to the reasoning performance.
In order to verify the effectiveness of thinking behavior alignment, we compare the average CoT draft length generated by the original Deepseek-R1-Distill-Qwen-1.5B and the fine-tuned version. Table 2 shows the mean CoT length on the 5 datasets. The CoT length generated by the fine-tuned model is significantly shorter than that of the original model on all datasets, especially for GSM8K. Therefore, thinking behavior alignment improves the efficiency of the drafting process.
4.4 Draft Selection Accuracy Analysis
In this section, we study the draft selection and error detection accuracy of the fine-tuned Deepseek-R1-Distill-Qwen-32B on the 500 samples in the test set of GSM8K. We divide these samples into two categories according to their labels. Samples for which the draft model can generate the correct reasoning chain form Class 1 (the label $y\neq n+1$), while others constitute Class 2 (the label $y=n+1$). Table LABEL:tab:selection shows the statistical results. There are 450 first-class samples out of the 500 samples. Since the second type of sample is more difficult to predict and the training data is limited, the target model demonstrates more accurate prediction for Class 1 samples than Class 2 samples. Our preliminary tests show that direct prediction by the original target model on the second-class samples yields zero accuracy. After fine-tuning, the model can correctly predict the labels for about half of the Class 2 samples.
4.5 Impact on Length of the Final Answer
We study the impact of SCoT on the length of the final answer in this section. Table 3 displays the answer length of the original 32B model and SCoT on the 5 datasets. In addition to reducing the latency of the reasoning phase, SCoT also reduces the length of the generated final answer compared to the target model and even brings additional inference efficiency improvements.
Table 3: Average length of the final answer generated by Deepseek-R1-Distill-Qwen-32B and SCoT.
Method GSM8K MATH GaoKao CollegeMath Olympiad
Deepseek-R1-Distill-Qwen-32B 314 632 948 416 2952
SCoT (32B+1.5B) 270 358 610 357 2179
5 Conclusion
In this paper, we proposed a novel reasoning framework called Speculative Chain-of-Thought to alleviate the latency bottleneck of reasoning LLMs. The proposed method leverages the efficiency of small models for parallel CoT drafting and the higher accuracy of the large model for drafting selection and error correction. The draft selection mechanism adaptively decides whether the target model needs to rethink based on the difficulty of the problem and the quality of the draft generation. Thought chains are mostly generated by the draft model for simple tasks. While more thought chains will be allocated to the target model to generate for complex problems. Speculative Chain-of-Thought demonstrates a speed-up ratio of up to 2.92× for the reasoning process while achieving near-target-model-level accuracy on 5 datasets of different difficulties. It exhibits the potential for future work to realize a better accuracy-efficiency trade-off for reasoning LLMs through model collaboration.
Appendix A Sample Presentation
We display two samples of constructed training data for draft selection in Figure 5 and Figure 6.
Figure 5: The selection template is in red and question is in blue.
Figure 6: The selection template is in red and question is in blue.